# dragapart.github.io

### v1

@@ Abstract
We introduce DragAPart, a method that, given an image and a set of drags as input, can generate a new image of the same object in a new state, compatible with the action of the drags. Differently from prior works that focused on repositioning objects, DragAPart predicts part-level interactions, such as opening and closing a drawer. We study this problem as a proxy for learning a generalist motion model, not restricted to a specific kinematic structure or object category. To this end, we start from a pre-trained image generator and fine-tune it on a new synthetic dataset, Drag-a-Move, which we introduce. Combined with a new encoding for the drags and dataset randomization, the new model generalizes well to real images and different categories. Compared to prior motion-controlled generators, we demonstrate much better part-level motion understanding.
@@
Figure 1: Examples of our DragAPart. Differently from prior works that used drags to move an object, each drag in DragAPart represents a part-level interaction, resulting in a physically plausible deformation of the object shape. DragAPart is trained on a new synthetic dataset, Drag-a-Move, for this task, and generalizes well to real data and even unseen categories. The trained model also can be used for segmenting movable parts and analyzing motion prompted by a drag.
@@ 1 Introduction
We consider the problem of learning an interactive image generator that allows to move the parts of an object by dragging them [63, 45], instead of repositioning the object as a whole. For example, dragging on the door of a cabinet should result in the image of the same cabinet, but with the door open (Fig. 1).
@@
Besides applications to controlled image generation, we explore dragging as a way of learning and probing generalist models of motion. Modelling deformable objects often uses ad-hoc models specific to each object type. For example, the SMPL [40] only represents humans and SMAL [79] only certain mammals. Some works [72, 26, 71, 35] have suggested to train similar models automatically from raw Internet images, but still assume that objects share the same kinematic structure (a skeleton), which limits generality. In stark contrast, foundation models like CLIP [50], GPT-4 [43], DALL-E [51] and Stable Diffusion (SD) [53], take a generalist approach. They do not pose any restriction on the type of content in the images or text. Likewise, a foundation model of motion should be able to understand the motion of any object, whether it is a human, a jellyfish, a towel, or a piece of furniture. However, there is no universal deformable template that can express the pose and deformation of all objects in a homogeneous manner — for example, the pose of a human and a piece of furniture are hardly comparable.
@@
We posit that a model of motion does not require to refer to a template; it is enough that the model understands the possible physical configurations of an object and their transitions. Dragging provides a way to probe such a model without using a template. A drag specifies how a single physical point of the object moves, leaving the model to ‘fill in’ the details by predicting a plausible motion for the object’s parts. For example, dragging a single drawer in a cabinet should not result in all other drawers moving together, because their motion is independent, but dragging the handle of a microwave oven’s door should open the door itself, as the motion is in this case perfectly correlated (Fig. 1).
@@
The dragging approach also lends itself to large-scale training, potentially from millions of Internet images and videos, by exploiting the strengths of recent generative models [51, 53, 60, 19]. In fact, it is natural to ask whether existing off-the-shelf image generators [51, 53] might already understand dragging and thus motion. Like others before us [59, 42, 63, 17], we found that these models can indeed respond to dragging, but only in the coarse sense of shifting an entire object, but fail to capture more nuanced effects. Hence, the challenge is to encourage image generative models to understand fine-grained and thus informative dragging outcomes.
@@
Our first contribution is to develop Drag-a-Move, a dataset to make generative models to understand part-level dragging. Previous works [58, 39, 33, 78] have proposed to use synthetic 3D objects [14, 13] to finetune pre-trained generative models for new view synthesis (NVS). Because the base model is trained on billions of images, a relatively small number of such objects, in the order of millions, is sufficient to train generalizable NVS. We thus propose to take a similar approach and contribute data rendered from an existing 3D dataset with rich part-level annotations [18]. Augmented with drag annotations, we can then use this data to finetune the generator and learn non-trivial drag responses. Even so, our synthetic data is far less abundant than that available for new view synthesis, which may include several million objects [14, 13]. Instead, our data is in the hundreds of different synthetic 3D objects due to the cost of annotating parts and kinematics. While this is not a large dataset, we show that texture randomization can significantly improve generalization to out-of-domain cases, including real data and unseen catagories at test time.
@@
We then use our dataset to train a new model, DragAPart, for the task of part-oriented drag control. Key to the design of this network is how the drags are encoded. Compared to prior works that considered similar tasks, we thus propose a new way of encoding drags with benefits that are transversal to the underlying model architecture. We also explored different generator architectures for this: DiT [46] with a transformer backbone trained from scratch on Drag-a-Move and Stable Diffusion [53] with a U-Net [54] backbone pretrained on billions of images, and study their performance on synthetic and real data.
@@
Finally, we explore some downstream applications of DragAPart. First, we show that it can be used to optimize the motion parameters of a given articulated 3D object to predict how its movable part is likely to move, subject to a drag. Second, we show that the model’s implicit part-level understanding can be leveraged to segment moving parts in an image, prompted by a drag.
@@ 2 Related Work
Generative models, such as diffusion models [22, 61, 62], can generate high-quality images [51, 53, 56], videos [21, 4] and 3D assets [64, 32, 37]. Conditioning plays a key role in controlling the generated contents, as well as its quality. For instance, DALL-E 3 [3] emphasise the importance of the quality of the textual prompt for training. ControlNet [76] show how to inject new kind of controls in a pre-trained diffusion model, and is at the basis of a new model zoo [48], widely used in applications such as text-to-image/video/3D [55, 20, 49] and image-to-image/video/3D [24, 78, 69, 26].
@@
Authors have considered controlling image and video generation via dragging. Training-free methods iteratively update the source image to match user-specified drags. Among those, DragGAN [45] optimizes a latent representation of the image in StyleGAN [30] to match the user-specified drags. DragDiffusion [59] ports this idea to SD [53]. DragonDiffusion [42] and Motion Guidance [17] combine SD with a guidance term [15] that capture feature correspondences and a flow loss. Training-based methods, on the other hand, learn drag-based control using ad-hoc training data for this task. For instance, iPoke [5] trains a variational autoencoder (VAE) to synthesize videos with objects in motion. MCDiff [10] and YODA [12] train a diffusion model using DDPM and flow matching, respectively. Li et al. [34] use a Fourier-based representation of motion suitable for natural, oscillatory dynamics characteristic of objects like trees and candles, and generates motion with a diffusion model. DragNUWA [75] and MotionCtrl [68] extend text-to-video generators with drag-based control.
Many of these methods show impressive results in term of repositioning objects in a scene, but do not allow to control motion at the level of object parts. Here, we address the latter task by introducing a new synthetic dataset and a new encoding technique for the drags.
@@ 3 Method
We develop DragAPart, an “interactive” generative model that, given as input a single object-centric RGB image $y$ and one or more drags $\mathcal{D}$, synthesizes a second image $x\sim\mathbb{P}(x|y,\mathcal{D})$ that reflects the effect of the drags. We first provide a formal definition of the problem and discuss the challenges of learning such a motion prior for articulated objects in Sec. 3.1. To overcome these challenges, we introduce a novel mechanism to encode the drags $\mathcal{D}$ (i.e., the sparse motion condition), which enables more efficient information propagation, in Sec. 3.2. Since we train our model only on synthetic renderings, in Sec. 3.3, we propose a simple yet effective way to mitigate the sim-to-real gap. In Sec. 3.4, we suggest downstream applications of DragAPart.
@@ 3.1 Preliminaries
The images $x,y\in\mathbb{R}^{3\times\Omega}$ are tensors defined on the spatial grid $\Omega=\{1,\dots,H\}\times\{1,\dots,W\}.$ A drag $d$ is a pair $(u,v)\in\Omega\times\mathbb{Z}^{2}$. The origin $u$ of the drag is always contained in the domain $\Omega$ of image $y$, but the termination $v$ may not, since the selected point may move outside the camera’s field of view. Given the image $y$ and one or more drags $\mathcal{D}\subset\Omega\times\mathbb{Z}^{2}$, the goal is to draw samples from the conditional distribution $\mathbb{P}(x|y,\mathcal{D})$.
@@
The epistemology of this model is defined by the expected effect of the drags. Each image contains a certain object, such as a cabinet or a microwave oven, and the effect of the drag is to interact with the corresponding object part. Formally, let $M\subset\mathbb{R}^{3}$ be the (unknown) surface of the 3D object contained in the image and consider a drag $(u,v)\in\Omega\times\mathbb{Z}^{2}$. Let $p\in M$ be the physical object point that corresponds to the drag origin $u$. Then the effect of the drag $(u,v)$ is to cause the object to deform such that point $p$ moves to a new location $q\in\mathbb{R}^{3}$ that re-projects as close as possible to the drag termination $v$, while obeying the physical rules that govern the object’s deformation. The effect of multiple drags is to specify multiple such constraints, which must be satisfied simultaneously. Naturally, there is no explicit model of the 3D shape or kinematic properties of the object in the image. Instead, the model must learn these properties implicitly.
@@
Note that not all drags are physically plausible, particularly if the drags are given by a user, so they may not be perfectly physically consistent with respect to the underlying 3D object’s geometry. Hence, the model must simulate a motion of the object that corresponds to the drags as well as possible, but without breaking physical plausibility. On the other hand, drags often under-specify the motion, which justifies using a stochastic rather than deterministic prediction model. In particular, a single drag can often be satisfied in a physically plausible manner by translating the object as a whole. This solution is not very informative about the object’s dynamics. Instead, we are interested in modeling nuanced physical interactions at the level of the individual object parts.
@@
Such nuanced interactions are typical in natural data. For example, dragging on the handle of a drawer normally results in opening the drawer instead of translating the entire cabinet, so, in principle, should be learned automatically from, e.g., raw Internet data. Perhaps surprisingly, however, we found that this is not necessarily the case. Following the lead of prior motion-conditioned image generators [45, 59, 42, 17], the obvious approach is to start from a pre-trained foundation model like Stable Diffusion [53], and then modify its internal representation of the image $y$ to respond to drags. The advantage is that foundation models are trained on billions of Internet images, and as such should have an understanding of almost any object and their configurations. However, we found that such methods tend to move the object as a whole at best, which is uninteresting for us. We hypothesize that the reason is that the training data of foundation models, while abundant in quantity, inevitably entangles multiple factors such as camera viewpoint and multi-object interactions, making it difficult to understand differnt part-level object configurations.
@@ 3.2 DragAPart: Architecture and Drag Encoding
Given the challenges identified in Sec. 3.1, we now describe DragAPart, which learns a motion prior for articulated objects from a pre-trained image generator. Key to our model is to fine-tune the motion generator, which is a distribution $\mathbb{P}(x|y,\mathcal{D})$, on a synthetic dataset of triplets $(x,y,\mathcal{D})$. This dataset will be described in Sec. 4. In addition to this data, which is not very large compared to datasets used to train image generators, we must also address the challenge of generalization. The rest of this section describes our architectural choices which improve the model’s generalizability, both across domains to real-world images and across categories to those that are not present in the dataset.
@@
We build our model on large pre-trained diffusion models for image generation, such as Stable Diffusion (SD) [53]. These models are trained on a very large number of natural images (i.e. LAION-5B [57]) and are open-ended. However, they are controlled by textual prompts, whereas our conditioning input consists of a reference image $y$ and a set of drags $\mathcal{D}$. Hence, our goal is to modify the model to accept the pair $(y,\mathcal{D})$ as condition.
@@
For the network $\Phi$, we experiment with two diffusion architectures, SD and DiT. The SD model is a UNet with several transformer blocks $l=1,2,\dots,L$ operating at different spatial resolutions $(h_{l},w_{l})$. It is conditioned on a textual prompt encoded using CLIP [50]. At each UNet block $l$, a self-attention layer generates spatial queries, keys and values $Q_{l},K_{l},V_{l}\in\mathbb{R}^{C_{l}\times h_{l}\times w_{l}}$ at resolution $(h_{l},w_{l})$. In order to inject the source image $y$ as a condition, we follow [8, 70] and encode $y$ utilizing the same SD encoder. Then, the keys and values $K_{l}^{y},V_{l}^{y}$ for image $y$ are used to replace the keys and values $K_{l},V_{l}$ for the generated image $x$, thus turning self-attention in cross-attention. In order to encode the drags $\mathcal{D}$, we consider a drag encoder $F_{l}(\mathcal{D})\in\mathcal{R}^{C_{l}\times h_{l}\times w_{l}}$ outputting an encoding of the drags at a resolution compatible with block $l$ in the transformer. These features are concatenated along the channel dimension to the input of each transformer block. The transformer layers are modified to handle the additional input channels, zero-initializing [76] the corresponding model weights.
The DiT model has several attention blocks at the same latent resolution. For each block, we follow [46] and inject the (tokenized) drag encoding via a token-wise adaLN-Zero block [47]. We defer the details to Sec. 0.A.1.
@@
We now define the encoding $F_{l}(\mathcal{D})$ of the drags. Prior works [75, 68] represent $\mathcal{D}$ as a sparse optical flow image $F(\mathcal{D})\in\mathbb{R}^{2\times\Omega}$ which is all zero except that, for each drag $(u,v)\in\mathcal{D}$, location $u$ contains the displacement $v-u$. This mechanism has two shortcomings. First, the transformer can sense well the origin $u$ of the drag, which is encoded spatially, but cannot do so for the termination $v$. Second, because all drags are encoded by a single optical flow image and the latter must be reduced to resolution $(h_{l},w_{l})$ before it can be injected in the LDM transformer, distinct drags may overlap. For instance, the works of [75, 68] first blur the sparse flow image $F$ using a Gaussian filter and then apply further convolutional layers to the result before injecting the information in the LDM. As demonstrated in Sec. 5, this design can confuse different drags, and diffuse a drag so that it affects more than the intended part (Fig. 5 left).
In order to mitigate these downsides, we introduce a new encoding $F_{l}$ for the drags, illustrated in Fig. 2. First, we assume a maximum number $|\mathcal{D}|\leq N$ of drags and assign different channels to different ones, so that they do not interfere with each other in the encoding. Second, we use separate channels to encode the drag source $u$ and the drag termination $v$. Specifically, given a single drag $d=(u,v)$ where $u=(u_{h},u_{w})$, let $F_{l}(u,v)\in\mathbb{R}^{2\times h_{l}\times w_{l}}$ be the tensor of all negative ones [^1] So we distinguish between no drag and drags precisely between two latent pixels. except for $F_{l}\left(\left\lfloor\frac{u_{h}\cdot h_{l}}{H}\right\rfloor,\left\lfloor\frac{u_{w}\cdot w_{l}}{W}\right\rfloor\right)=\left(\frac{u_{h}\cdot h_{l}}{H}-\left\lfloor\frac{u_{h}\cdot h_{l}}{H}\right\rfloor,\frac{u_{w}\cdot w_{l}}{W}-\left\lfloor\frac{u_{w}\cdot w_{l}}{W}\right\rfloor\right)\in[0,1]^{2}$. We then define $F_{l}(\mathcal{D})=\bigoplus_{(u,v)\in\mathcal{D}}F_{l}(u,v)\oplus F_{l}(v,u)\in\mathcal{R}^{4N\times h\times w}$ to be the concatenation of the encoding of all drags in $\mathcal{D}$ along the channel dimension. Note that we encode each drag twice, swapping source and termination. [^2] There is a caveat when the drag termination $v$ is outside the image, in which case $\left(\left\lfloor\frac{u_{h}\cdot h}{H}\right\rfloor,\left\lfloor\frac{u_{w}\cdot w}{W}\right\rfloor\right)$ is outside resolution $(h,w)$, so we define the termination $v$ at the nearest spatial location inside $(h,w)$. The corresponding values will also be relative to this new coordinate, so they might not lie in $\left[0,1\right]$ (see Fig. 2 for an example). If there are less than $N$ drags, we simply pad the tensor with zero. We show in Sec. 5 that our proposed multi-resolution drag encoding technique boosts the generation quality in both U-Net and transformer-based LDMs.

### v2

@@ Abstract
We introduce DragAPart, a method that, given an image and a set of drags as input, generates a new image of the same object that responds to the action of the drags. Differently from prior works that focused on repositioning objects, DragAPart predicts part-level interactions, such as opening and closing a drawer. We study this problem as a proxy for learning a generalist motion model, not restricted to a specific kinematic structure or object category. We start from a pre-trained image generator and fine-tune it on a new synthetic dataset, Drag-a-Move, which we introduce. Combined with a new encoding for the drags and dataset randomization, the model generalizes well to real images and different categories. Compared to prior motion-controlled generators, we demonstrate much better part-level motion understanding.
@@
Figure 1: Examples of our DragAPart. Each drag in DragAPart represents a part-level interaction, resulting in a physically plausible deformation of the object shape. DragAPart is trained on a new synthetic dataset, Drag-a-Move, for this task, and generalizes well to real data and even unseen categories. The trained model also can be used for segmenting movable parts and analyzing motion prompted by a drag.
@@ 1 Introduction
We consider the problem of learning an interactive image generator that allows moving the parts of an object by dragging them [61, 43] instead of repositioning the object as a whole. For example, dragging on the door of a cabinet should result in the image of the same cabinet but with the door open (Fig. 1).
@@
Besides applications to controlled image generation, we explore dragging as a way of learning and probing generalist models of motion. Modelling deformable objects often uses ad-hoc models specific to each object type. For example, SMPL [38] only represents humans and SMAL [75] only certain mammals. Some works [69, 24, 68, 33] have suggested training similar models automatically from raw Internet images but still assuming that objects share the same kinematic structure (a skeleton) which limits generality. In stark contrast, foundation models like CLIP [48], GPT-4 [41], DALL-E [49] and Stable Diffusion (SD) [51] take a generalist approach. They do not pose any restriction on the type of content in the images or text. Likewise, a foundation model of motion should be able to understand the motion of any object, whether it is a human, a jellyfish, a towel, or a piece of furniture. However, there is no universal deformable template that can express the pose and deformation of all objects in a homogeneous manner — for example, the pose of a human is hardly comparable to that of a cabinet.
@@
We posit that a model of motion does not require to refer to a template; it is enough that the model understands the possible physical configurations of an object and their transitions. Dragging provides a way to probe such a model without using a template. A drag specifies how a single physical point of the object moves, leaving the model to ‘fill in’ the details by predicting a plausible motion for the object’s parts. For example, dragging a single drawer in a cabinet should not result in all other drawers moving together because their motion is independent, but dragging the handle of a microwave oven’s door should open the door itself, as the motion is, in this case, perfectly correlated (Fig. 1).
@@
The dragging approach also lends itself to large-scale training, potentially from millions of Internet images and videos, by exploiting the strengths of recent generative models [49, 51, 58, 18]. In fact, it is natural to ask whether existing off-the-shelf image generators [49, 51] might already understand dragging and thus motion. Like others [57, 40, 61, 16], we found that these models do respond to dragging, but only in the sense of shifting an entire object, but fail to capture more nuanced effects. Hence, the challenge is to encourage generative models to produce fine-grained and, thus informative dragging outcomes.
@@
Our first contribution is to develop Drag-a-Move, a dataset to align image generators with part-level dragging. Previous works [56, 37, 31, 74] have proposed to use synthetic 3D objects [14, 13] to fine-tune pre-trained generative models for new view synthesis (NVS). Because the base model is trained on billions of images, a relatively small number of such objects, in the order of millions, is sufficient to train generalizable NVS. We thus propose to take a similar approach and contribute data rendered from an existing 3D dataset with rich part-level annotations [17]. Augmented with drag annotations, this data can be used to fine-tune the generator and learn non-trivial drag responses. Even so, our synthetic data is far less abundant than that available for NVS, which includes several million objects [14, 13]. Instead, our data is in the hundreds of different synthetic 3D objects due to the cost of annotating parts and kinematics. While the number of objects is small, we show that texture randomization can significantly improve generalization to out-of-domain cases, including real data and unseen categories at test time.
@@
We then use our dataset Drag-a-Move to train a new model, DragAPart, for the task of part-oriented drag control. Key to the design of DragAPart is how the drags are encoded. Compared to prior works that considered similar tasks, we propose a new way of encoding drags with benefits that are transversal to the underlying model architecture. We explore different network architectures for this task: DiT [44] with a transformer [64] backbone trained from scratch and Latent Diffusion architecture [51] with a U-Net [52] backbone pre-trained on billions of images, and study their performance on synthetic and real data.
@@
Finally, we also explore some downstream applications of DragAPart. First, we show that it can be used to optimize the motion parameters of a given articulated 3D object to predict how its movable part is likely to move, subject to a drag. Second, we show that the model’s implicit part-level understanding can be leveraged to segment moving parts in an image prompted by a drag.
@@ 2 Related Work
Generative models, such as diffusion models [21, 59, 60], can generate high-quality images [49, 51, 54], videos [20, 4] and 3D assets [62, 30, 35]. Conditioning plays a key role in controlling the generated content and its quality. For instance, DALL-E 3 [3] emphasizes the importance of quality textual prompts used for training. ControlNet [73] shows how to inject new kinds of controls in a pre-trained diffusion model and is at the basis of a new model zoo [46], widely used in applications such as text-to-image/video/3D [53, 19, 47] and image-to-image/video/3D [22, 74, 66, 24].
@@
Authors have considered controlling image and video generation via dragging. Training-free methods iteratively update the source image to match user-specified drags. Among those, DragGAN [43] optimizes a latent representation of the image in StyleGAN [28] to match the user-specified drags. DragDiffusion [57] ports this idea to SD [51]. DragonDiffusion [40] and Motion Guidance [16] combine SD with a guidance term that captures feature correspondences and a flow loss. Training-based methods, on the other hand, learn drag-based control using ad-hoc training data for this task. For instance, iPoke [5] trains a variational autoencoder (VAE) to synthesize videos with objects in motion. MCDiff [10] and YODA [12] train a diffusion model using DDPM and flow matching, respectively. Li et al. [32] use a Fourier-based representation of motion suitable for natural, oscillatory dynamics characteristic of objects like trees and candles and generates motion with a diffusion model. DragNUWA [72] and MotionCtrl [65] extend text-to-video generators with drag-based control.
Many of these methods show impressive results in terms of repositioning objects in a scene, but do not allow to control motion at the level of object parts. Here, we address the latter task by introducing a new synthetic dataset and a new encoding technique for the drags.
@@ 3 Method
We develop DragAPart, an interactive generative model that, given as input a single object-centric RGB image $y$ and one or more drags $\mathcal{D}$, synthesizes a second image $x\sim\mathbb{P}(x|y,\mathcal{D})$ that reflects the effect of the drags. We first provide a formal definition of the problem and discuss the challenges of learning such a motion prior for articulated objects in Sec. 3.1. To overcome these challenges, we introduce a novel mechanism to encode the drags $\mathcal{D}$ (i.e., the sparse motion condition), which enables more efficient information propagation in Sec. 3.2. Since we train our model only on synthetic renderings, in Sec. 3.3, we propose a simple yet effective way to mitigate the sim-to-real gap. In Sec. 3.4, we suggest downstream applications of DragAPart.
@@ 3.1 Preliminaries
We consider images $x,y\in\mathbb{R}^{3\times\Omega}$ as tensors defined on the spatial grid $\Omega=\{1,\dots,H\}\times\{1,\dots,W\}.$ A drag $d$ is a pair $(u,v)\in\Omega\times\mathbb{Z}^{2}$. Given the image $y$ and one or more drags $\mathcal{D}\subset\Omega\times\mathbb{Z}^{2}$, the goal is to draw samples from the conditional distribution $\mathbb{P}(x|y,\mathcal{D})$. The origin $u$ of the drag is always contained in the domain $\Omega$ of image $y$, but the termination $v$ may not be since the selected point may move outside the camera’s field of view.
@@
The distribution $\mathbb{P}(x|y,\mathcal{D})$ is defined by the expected effect of the drags. Each image contains a certain object, such as a cabinet or a microwave oven, and the effect of each drag is to interact with the corresponding object part. Formally, let $M\subset\mathbb{R}^{3}$ be the (unknown) surface of the 3D object contained in the image and consider a drag $(u,v)\in\Omega\times\mathbb{Z}^{2}$. Let $p\in M$ be the physical object point that corresponds to the drag origin $u$. Then, the effect of the drag $(u,v)$ is to cause the object to deform such that point $p$ moves to a new location $q\in\mathbb{R}^{3}$ that re-projects as close as possible to the drag termination $v$, while obeying the physical rules that govern the object’s deformation. The effect of multiple drags is to specify multiple such constraints, which must be satisfied simultaneously. Because there is no explicit model of the 3D shape or kinematic properties of the object in the image, the model must learn these properties implicitly.
@@
Note that not all drags are physically plausible, particularly if the drags are given by a user, so they may not be perfectly physically consistent with respect to the underlying 3D object’s geometry. Hence, the model must simulate a motion of the object that corresponds to the drags as well as possible, but without breaking physical plausibility. On the other hand, drags often under-specify the motion, which justifies using a stochastic rather than deterministic prediction model. In particular, a single drag can often be satisfied in a physically plausible manner by translating the object as a whole. This solution is not very informative about the object’s dynamics. Instead, we are interested in modelling nuanced physical interactions at the level of the individual object parts.
@@
Such nuanced interactions are typical in natural data — for example, dragging on the handle of a drawer normally results in opening the drawer instead of translating the entire cabinet — so, in principle, should be learned automatically from, e.g., raw Internet data. Perhaps surprisingly, however, we found that this is not necessarily the case. Following the lead of prior motion-conditioned image generators [43, 57, 40, 16], the obvious approach is to start from a pre-trained foundation model like Stable Diffusion [51], and then modify its internal representation of the image $y$ to respond to drags. The advantage is that foundation models are trained on billions of Internet images and, as such, should have an understanding of almost any object and its configurations. However, we found that such methods tend to move the object as a whole at best, which is uninteresting for us. We hypothesize that the reason is that the training data of foundation models, while abundant in quantity, inevitably entangles multiple factors such as camera viewpoint and multi-object interaction, making it difficult to understand different part-level object configurations.
@@ 3.2 DragAPart: Architecture and Drag Encoding
Given the challenges identified in Sec. 3.1, we now describe DragAPart, which learns a motion prior for articulated objects from a pre-trained image generator. Key to our model is to fine-tune the motion generator, which models the distribution $\mathbb{P}(x|y,\mathcal{D})$, on a synthetic dataset of triplets $(x,y,\mathcal{D})$. This dataset will be described in Sec. 4. In addition to this data, which is not very large compared to datasets used to train image generators, we must also address the challenge of generalization. The rest of this section describes our architectural choices which improve the model’s generalizability, both across domains to real-world images and across categories to those that are not present in the dataset.
@@
We build our model on large pre-trained diffusion models for image generation, such as Stable Diffusion (SD) [51]. These models are trained on a very large number of natural images (i.e. LAION [55]) and are open-ended. However, they are controlled by textual prompts, whereas our conditioning input consists of a reference image $y$ and a set of drags $\mathcal{D}$. Hence, our goal is to modify the model to accept the pair $(y,\mathcal{D})$ as conditions.
@@
For the network $\Phi$, we experiment with two diffusion architectures, SD and DiT. The SD model is a UNet with several transformer blocks $l=1,2,\dots,L$ operating at different spatial resolutions $(h_{l},w_{l})$. It is conditioned on a textual prompt encoded using CLIP [48]. At each UNet block $l$, a self-attention layer generates spatial queries, keys and values $Q_{l},K_{l},V_{l}\in\mathbb{R}^{C_{l}\times h_{l}\times w_{l}}$ at resolution $(h_{l},w_{l})$. In order to inject the source image $y$ as a condition, we follow [8, 67] and pass $y$ (without adding noise to it) through the same network $\Phi$. Then, the keys and values $K_{l}^{y},V_{l}^{y}$ from image $y$ are used to replace the keys and values $K_{l},V_{l}$ for the generated image $x$, thus turning self-attention into cross attention between the noised image and the clean reference image $y$. In order to encode the drags $\mathcal{D}$, we consider a drag encoder $F_{l}(\mathcal{D})\in\mathcal{R}^{C_{l}\times h_{l}\times w_{l}}$ outputting an encoding of the drags at the resolution compatible with transformer block $l$. These features are concatenated along the channel dimension to the input of each transformer block. The transformer layers are modified to handle the additional input channels, zero-initializing the corresponding model weights.
The DiT model has several attention blocks at the same latent resolution. For each block, we follow [44] and inject the (tokenized) drag encoding via a token-wise adaLN-zero block [45]. We defer the details to the sup. mat.
@@
We now define the encoding $F_{l}(\mathcal{D})$ of the drags. Prior works [72, 65] represent $\mathcal{D}$ as a sparse optical flow image $F(\mathcal{D})\in\mathbb{R}^{2\times\Omega}$ which is all zero except that, for each drag $(u,v)\in\mathcal{D}$, location $u$ contains the displacement $v-u$. This mechanism has two shortcomings. First, the transformer can sense well the origin $u$ of the drag, which is encoded spatially, but cannot do so for the termination $v$. Second, because all drags are encoded by a single optical flow image and the latter must be reduced to resolution $(h_{l},w_{l})$ before it can be injected into the LDM transformer block $l$, distinct drags may overlap. For instance, [72, 65] first blur the sparse flow image $F$ using a Gaussian filter and then apply further convolutional layers to the result before injecting the information into the LDM. As demonstrated in Sec. 5, this design can confuse different drags and diffuse a drag so that it affects more than the intended part (Fig. 5 left).
In order to mitigate these downsides, we introduce a new encoding $F_{l}$ for the drags, illustrated in Fig. 2. First, we assume a maximum number $|\mathcal{D}|\leq N$ of drags and assign different channels to different ones, so that they do not interfere with each other in the encoding. Second, we use separate channels to encode the drag source $u$ and the drag termination $v$. Specifically, given a single drag $d=(u,v)$ where $u=(u_{h},u_{w})$, let $F_{l}(u,v)\in\mathbb{R}^{2\times h_{l}\times w_{l}}$ be the tensor of all negative ones [^1] So we distinguish between no drag and drags precisely between two latent pixels. except for $F_{l}\left(\left\lfloor\frac{u_{h}\cdot h_{l}}{H}\right\rfloor,\left\lfloor\frac{u_{w}\cdot w_{l}}{W}\right\rfloor\right)=\left(\frac{u_{h}\cdot h_{l}}{H}-\left\lfloor\frac{u_{h}\cdot h_{l}}{H}\right\rfloor,\frac{u_{w}\cdot w_{l}}{W}-\left\lfloor\frac{u_{w}\cdot w_{l}}{W}\right\rfloor\right)\in[0,1]^{2}$. We then define $F_{l}(\mathcal{D})=\bigoplus_{(u,v)\in\mathcal{D}}F_{l}(u,v)\oplus F_{l}(v,u)\in\mathcal{R}^{4N\times h\times w}$ to be the concatenation of the encoding of all drags in $\mathcal{D}$ along the channel dimension. Note that we encode each drag twice, swapping source and termination. [^2] There is a caveat when the drag termination $v$ is outside the image, in which case $\left(\left\lfloor\frac{u_{h}\cdot h}{H}\right\rfloor,\left\lfloor\frac{u_{w}\cdot w}{W}\right\rfloor\right)$ is outside resolution $(h,w)$, so we define the termination $v$ at the nearest spatial location inside $(h,w)$. The corresponding values will also be relative to this new coordinate, so they might not lie in $\left[0,1\right]$ (see Fig. 2 for an example). If there are less than $N$ drags, we simply pad the tensor with $-1$. We show in Sec. 5 that our proposed multi-resolution drag encoding technique boosts the generation quality in both U-Net and attention-based LDMs.